# Extension principles for the Einstein Yang–Mills system

### v1

@@ Abstract.
We establish an extension principle for the spherically symmetric Einstein Yang–Mills system (SSEYM) with $H^{1}$ data. Based on this result, we further prove an extension theorem for developments of weighted $H^{1}$ data. In particular, the weighted $H^{1}$ space allows Hölder continuous data. Therefore, our result is consistence with the conjecture that well-posedness holds for Hölder continuous data (with sufficiently many angular derivatives) in vacuum where the axis is included.
In contrast to a massless scalar field, the purely magnetic Yang–Mills field in spherical symmetry satisfies a wave-type equation with singular potential. The proof of Christodoulou [10], based on an $L^{\infty}-L^{\infty}$ estimate, fails for the Yang–Mills scenario. Instead, we employ the $L^{2}$ method, which works for the (massless or massive) scalar matter field as well.
@@ 1. Introduction
In a series of works [9, 10, 11, 13], Christodoulou proved the weak cosmic censorship conjecture [^1] Generic asymptotically flat initial data have a maximal future development possessing a complete future null infinity[12]. for the spherically symmetric Einstein massless scalar field system. It was further demonstrated by Li-Liu [16, 17] and An [1] under gravitational perturbations without symmetries. Recently, An-Tan [2] gave a proof for the system of Einstein charged massless scalar field in spherical symmetry as well.
The instability mechanic of naked singularity in [13] was found out in the topology of bounded variation (BV). From the other perspective, the local existence theory and further the small data global existence in the BV topology were both verified ahead [10]. As shown in [10, 13], the proofs of these (local or global) existence or instability statements were all built on a sharp criterion for $C^{1}$ extension. Along the line of [10], extension criterions were verified as well in the context of Einstein charged massless scalar field and Einstein massless Vlasov system in spherical symmetry [2, 20].
@@
We are now concerned with the Einstein Yang–Mills system, which is significant for various reasons. From a physical perspective, it triggers a wealth of fascinated gravitational collapses, as seen, for instance, [3, 6, 5] for several numerical results. On the other hand, we will show in this paper that the mathematical techniques in the treatment of Einstein Yang–Mills system will be significantly different from those for gravity coupled with massless matter fields [10, 20, 2].
@@
In this paper, we initiate the study of extension principle for the Einstein Yang–Mills system in spherical symmetry. We will first consider the characteristic problem for the spherically symmetric Einstein Yang–Mills (SSEYM) system where initial data are prescribed on a future light cone emanating from vertex at the center of symmetry. In this setting, we demonstrate the local existence theorem for data with a low-regularity norm ($H^{1}$ initial data). We also deduce an extension principle for developments of $H^{1}$ data. Our second result is an extension theorem in a weaker topology. We show that if the interior is vacuum (and hence Minkowski), the spacetime is extendable in the topology of weighted $H^{1}$ space. Following theses extension theorems, the Cauchy stabilities are justified for the Einstein Yang–Mills system under spherically symmetric perturbations, which are initially small with respect to the above norms respectively.
The weighted $H^{1}$ space allows the class of Hölder continuous data, or more precisely, data that are smooth except at one point and Hölder continuous there. It was conjectured by Rodnianski–Shlapentokh-Rothman [21] that well-posedness holds for Hölder continuous data with sufficiently many angular derivatives in vacuum (without symmetries) where the axis is included. Our extension theorem thus supports this conjecture. Previously, in the context of spherically symmetric massless matter field [10, 2], this conjecture is true in BV topology, where the instabilities of naked singularities are stated. However, as will be discussed later, the method used in massless case fails for Einstein–Yang–Mills scenario and hence a well-posed result in BV topology is not expected. The new $L^{2}$-based method we propose in this paper has a closer nature to the vacuum context, and may be relevant for future progress in vacuum.
@@ 1.1. Statement of the main results
We formulate the SSEYM system with the gauge group $SU(2)$ in Section 2.1, where we assume the spherically symmetric purely magnetic ansatz for the Yang–Mills potential (2.2). The characteristic problem involves prescribing a $C^{0}$ function $\partial_{r}w|_{C_{u_{0}}}$ on an initial outgoing cone $C_{u_{0}}$, where $w$ is related to the Yang–Mills potential (2.2). Throughout this paper, the $C^{0}$ Yang–Mills solution always refers to $\partial_{r}w\in C^{0}$.
@@
It suggests choosing the initial datum $w|_{C_{u_{0}}}$ can be replaced by prescribing $\partial_{r}w|_{C_{u_{0}}}$.
Let us begin our statement with the $H^{1}$ case.
1.1.1. Statement for the $H^{1}$ case
Fixing a constant
$0\leq b<\frac{\sqrt{89}-9}{4},$ (1.4)
@@ 1.1.1. Statement for the $H^{1}$ case
$\|\partial_{r}w\|_{H^{1}}:=\int_{C_{u}}\frac{1}{r^{1+b}}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}+(r\partial_{r}(\partial_{r}w))^{2}\right)\mathrm{d}r.$ (1.5)
@@
$\int_{0}^{r}\frac{1}{r^{1+b}}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r.$ (1.6)
@@
$\int_{0}^{r_{1}}\frac{1}{r^{1+b}}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r=\int_{-\infty}^{\ln r_{1}}e^{-bs}\sum_{0\leq k\leq 2}\left(\partial_{s}^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}s.$ (1.7)
When $b=0$, (1.7) is the standard $H^{2}$ norm for $\frac{\hat{w}}{r}$. We abuse the notation a little bit and call (1.7) the $H^{2}$ norm of $\frac{\hat{w}}{r}$, or rather $H^{1}$ norm of $\partial_{r}w$. With respect to this $H^{1}$ space, the density theorem is proved in Appendix C.
@@
As a further remark, the standard energy for the Yang–Mills field does not account for the definition of energy (1.5), even if $b=0$. Actually, the standard first order energy for the Yang–Mills field should be roughly in the form of $\int_{C_{u}}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}\right)\mathrm{d}r$, where the weight $\frac{1}{r^{1+b}}$ is absent. The motivation for (1.5) indeed comes from the right hand side of (2.14a), since after integrating (2.14a) in $\underline{u}$, we expect to bound $\frac{m}{r}$ via the $H^{1}$ (in fact $H^{0}$) norm of $\partial_{r}w$.
We present the $H^{1}$-extension theorem. One may refer to Theorem 3.1 for a comprehensive statement.
@@
Suppose we have the initial data $\partial_{r}w|_{C_{u_{0}}}$ with bounded $H^{1}$ norm, and an $H^{1}$ solution (that is, $\partial_{r}w\in H^{1}$) to the SSEYM system on $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$. There is a constant $\varepsilon>0$ such that if
@@
the solution extends as an $H^{1}$ solution on $\mathcal{D}(u_{0};\,\underline{u}_{1})$ for some $\underline{u}_{1}>\underline{u}_{\ast}$.
Following the extension theorem, there is the Cauchy stability statement, for which a more precise version is given in Theorem 3.9.
@@
Suppose the data $\partial_{r}w|_{C_{u_{0}}}$ admit bounded $H^{1}$ norm that is small enough. Then there is a global and unique solution on $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{0})$ with small $H^{1}$ norm, for any $\underline{u}_{0}\in(u_{0},+\infty)$ and $u_{1}\in(u_{0},\,\underline{u}_{0}]$.
We next proceed to the weighted $H^{1}$ case.
1.1.2. Statement for the weighted $H^{1}$ case
We will consider a class of weighted energies in the domain $\mathcal{D}(u_{0},u;\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$ and $\delta>0$. The weighted energies are defined with the assistance of a non-negative function
$\xi:=\frac{r-r_{\ast}}{r}\quad\text{with}\quad r_{\ast}(u):=r(u,\,\underline{u}_{\ast}),$ (1.8)
for $\underline{u}\geq\underline{u}_{\ast}$. We fix a constant $\gamma\in(b,2]$ with $b$ given in (1.4) and define the following weighted $H^{1}$ norm for $\partial_{r}w$
$\int_{C_{u}\cap\{\underline{u}\geq\underline{u}_{\ast}\}}\frac{1}{r^{1+b}}\left(\xi^{-3-\gamma}\frac{\hat{w}^{2}}{r^{2}}+\xi^{-1-\gamma}(\partial_{r}w)^{2}+\xi^{1-\gamma}\left(r\partial_{r}(\partial_{r}w)\right)^{2}\right)\mathrm{d}r.$ (1.9)
@@ 1.1.2. Statement for the weighted $H^{1}$ case
The weighted $H^{1}$ norm (1.9) is equivalent to
$\int_{\{r_{\ast}\leq r\}}\xi^{-3-\gamma}r^{-1-b}\sum_{0\leq k\leq 2}\left((\xi r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r|_{C_{u}}.$ (1.10)
For the density theorem in the weighted $H^{1}$ space, please refer to Appendix C for a proof.
Theorem 1.7.
Given datum $\partial_{r}w|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$ (namely, $\partial_{r}w|_{C_{u_{0}}\cap\{\underline{u}\leq\underline{u}_{\ast}\}}\equiv 0$), we assume $|r-r_{\ast}|^{-\frac{\gamma}{2}}\partial_{r}w|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$ and the weighted $H^{1}$ norm of $\partial_{r}w|_{C_{u_{0}}}$ on $C_{u_{0}}\cap\{\underline{u}_{\ast}\leq\underline{u}\leq\underline{u}_{\ast}+\delta\}$ for some $\delta>0$ is bounded. If $\delta$ is small enough, the solution extends as a solution with bounded, weighted $H^{1}$ norm in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$, where $u\in[u_{0},\,\underline{u}_{\ast})$.
Remark 1.8.
We remark that our data requirement allows the initially second derivative $\partial^{2}_{r}w|_{C_{u_{0}}}$ blows up [^2] The $C^{1}$ case ($\gamma=2$) is also included. like
$(r-r_{\ast})^{-1+\frac{\gamma}{2}},\quad 0\leq b<\gamma\leq 2.$
Thus, the data in Theorem 1.7 are, in fact, Hölder continuous. Roughly speaking, Theorem 1.7 states that if the interior $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is Minkowski, and the data are Hölder continuous, then the solution can be extended beyond the last incoming cone $\underline{C}_{\underline{u}_{\ast}}$. Furthermore, we will show that the extended solution is still Hölder continuous when it goes across $\underline{C}_{\underline{u}_{\ast}}$. Please refer to Theorem 4.1 for a detailed description.
The next one is a Cauchy stability statement in the weighted norm space. One may refer to Theorem 4.10 for more details.
Theorem 1.9.
Given datum $\partial_{r}w|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$, suppose $|r-r_{\ast}|^{-\frac{\gamma}{2}}\partial_{r}w|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$, and the weighted $H^{1}$ norm on $C_{u_{0}}\cap\{\underline{u}_{\ast}\leq\underline{u}\leq\underline{u}_{1}\}$, $\underline{u}_{1}<+\infty$, is small enough. Then there is a unique solution with bounded, weighted $H^{1}$ norm on $\mathcal{D}(u_{0},\,u;\,\underline{u}_{1})$, where $u\in[u_{0},\,\underline{u}_{\ast})$.
Remark 1.10.
In the subsection followed, we will sketch the main idea for the proofs. We propose the $L^{2}$ method, which is tailed for the Yang–Mills field, or rather, the wave type equation with a positive, singular potential. Indeed we can also adapt the $L^{2}$ scheme to a massless or massive scalar field so that all the preceding statements (Theorem 1.41.9) hold for the Einstein (massive or massless) scalar field system in spherical symmetry. Since our proof is independ of $C^{1}$ norm, but only depends on $H^{1}$ norm, we have in fact achieved an extension principle in a weaker topology ($H^{1}$ topology) than that in Christodoulou [10] ($C^{1}$ topology), and thus less regularity on the axis is obtained as well.
To facilitate the extension theorems, we verify the associated local existence theory for the characteristic problem of SSEYM system in Appendix A. We also state an extension principle away from the axis in Appendix B for completeness. Thus, it confirms that the “first singularity” for the SSEYM system must emanate from the axis.
@@ 1.2. Comments on the proof
The purely magnetic Yang–Mills field obeys a wave-type equation with a singular potential $\frac{1}{r}$, which makes itself significantly different from a standard massless scalar field. To highlight the difference, it is useful to compare their formulations in Minkowski space. Notice that, the massless sclar field equation in Minkowski takes the form of
@@
while the Yang–Mills equation reads
@@
In terms of $\hat{w}$ (1.3), the Yang–Mills equation can be reformulated as follows
@@
It manifests itself as a nonlinear wave equation with an additional linear, singular term
@@
compared to the scalar wave equation. As a consequence, the techniques for proving the extension principles are necessarily different.
@@ 1.2.1. The $H^{1}$ case
Remarkably, this equation contains no linear terms on the right hand side. Namely, terms that are linear on $\partial_{r}\alpha_{s}$, $\alpha_{s}$, $r\partial_{r}\phi$ and $\frac{2m}{r}$ are absent. It is exactly this feature that allows an usage of $L^{\infty}-L^{\infty}$ estimate in proving the extension theorem [10, Theorem 5.1]. Without loss of generality, we take the first term $-2\frac{\partial_{u}r}{r}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}\partial_{r}\alpha_{s}$ for instance to explain the main ingredient of this approach. In this quadratic term, the $\frac{2m}{r}$ offers smallness $\varepsilon$ so that the whole term is negligible and a type of inequality
@@
yields the bound for $\partial_{r}\alpha_{s}$. In principle, this kind of $L^{\infty}-L^{\infty}$ estimate works well for ”generic” massless matter fields that share the common feature of vanishing linear terms. Along with the massless scalar field, as far as we know, the $L^{\infty}-L^{\infty}$ strategy truns out to be valid as well for the charged massless scalar field [2] and the massless Vlasov system [20].
In contrast, for the Yang–Mills scenario, due to the additional potential, the main equation for $\partial_{r}\alpha$ (2.3), $\alpha:=\partial_{r}w$, involves linear terms that are linear on $\frac{\hat{w}}{r}$ and $\alpha$ respectively on the right hand side. Therefore, the strategy of $L^{\infty}-L^{\infty}$ estimate fails in the Yang–Mills case, since we are unable to gain extra smallness from these linear terms. Nevertheless, noticing the positive sign of the potential in the Yang–Mills equation, we propose the $L^{2}$ method, which allows us to treat the quadratic terms originated from the linear terms through integration by parts. After that, terms that have no additional smallness give rise to quadratic terms with definite signs. Since our energy is not the standard energy, it is in fact $r^{-1-b}$ weighted, these quadratic terms do not necessarily admit favorable signs. Specifically, we are unable to derive any energy bound using the first-order energy identities; see (3.1.4) and (3.1.4). To proceed, the idea is, after performing integration by parts on these linear-related terms, the second order energy identity generates positive or negative quadratic terms that are potential to compensate the corresponding terms with opposite signs in the first order energy identities, see (3.1.4). Then we play with these two energy identities by assigning different weights on them to render all the quadratic spacetime integrals positive. Eventually, we obtain the $H^{1}$ energy bounds, see Section 3.1.4, and the pointwise estimates follow from a Sobolev type inequality in the $H^{1}$ space.
The scheme employed for the Yang–Mills case applies to the (massless or massive) scalar field so that an extension principle in a low regularity ($H^{1}$) can be established. Moreover, using our method, we establish the extension principle with less regularity on the axis: $r\partial_{r}\phi\sim r^{b}$ with $0\leq b<\frac{\sqrt{89}-9}{4}$, while in Christodoulou [10], $r\partial_{r}\phi$ is roughly $r$ near the axis. Besides, the $H^{1}$ energy estimate is crucial to promote a local existence theorem, Theorem A.1.
1.2.2. The weighted $H^{1}$ case
The proof in the weighted $H^{1}$ case heavily depends on the energy identities derived in the previous $H^{1}$ case and its modification through introducing a weight function $\xi$ (1.8). Remarkably, the weight function $\xi$ is designed in such a way that in the domain concerned,
$0\leq\xi<1,\quad\partial_{u}\xi\geq 0,\quad\partial_{\underline{u}}\xi\geq 0$
More specifically, noting the negative power on $\xi$ in the first order weighted energy, the sign of $\partial_{u}\xi$ and $\partial_{\underline{u}}\xi$ helps to produce positive spacetime integrals which will cancel the negative terms in the originally unweighted energy identities. Once the first order estimate is completed, the second order one follows.

### v2

@@ Abstract.
We prove the local existence theorem and establish an extension principle for the spherically symmetric Einstein Yang–Mills system (SSEYM) with $H^{1}$ data. This in addition implies Cauchy stability for the system.
In contrast to a massless scalar field, the purely magnetic Yang–Mills field in spherical symmetry satisfies a wave-type equation with a singular potential. As a consequence, the proof of Christodoulou [10], based on an $L^{\infty}-L^{\infty}$ estimate, fails in the Yang–Mills context. Instead, we employ an $L^{2}$-based method, which is valid for both massless and massive scalar matter fields as well.
@@ 1. Introduction
In a series of works [9, 10, 11, 13], Christodoulou established a proof of the weak cosmic censorship conjecture [^1] Generic asymptotically flat initial data have a maximal future development possessing a complete future null infinity[12]. for the spherically symmetric Einstein massless scalar field system. This result was further demonstrated by Li-Liu [16, 17] and An [1] under gravitational perturbations without symmetries. More recently, An-Tan [2] also verified the weak cosmic censorship conjecture for the Einstein charged massless scalar field system under spherical symmetry.
As a core step in the proof, Christodoulou [13] introduced an instability mechanism for naked singularities within the bounded variation (BV) topology. This framwork is well motivated, since both the local existence theory and the subsequent small-data global existence result for the spherically symmetric Einstein massless scalar field system had already been established in the BV topology [10]. In detailed proofs, one observes that the arguments for both (local or global) existence and instability rely crucially on a sharp criterion for $C^{1}$ extension [10, 13]. This indicates that an extension principle in a suitable topology serves as the foundational element for addressing the weak cosmic censorship conjecture. Following the methodology of [10], such extension criteria have also been established in the context of the Einstein charged massless scalar field and the Einstein massless Vlasov system under spherical symmetry[2, 20].
@@
We are now concerned with the Einstein Yang–Mills system, which is of interest for several reasons. From a physical perspective, it triggers a rich variety of fascinated gravitational collapses, as evidenced by numerous numerical studies (see, for instance, [3, 6, 5]). On the mathematical side, we demonstrate in this paper that the techniques required for the Einstein Yang–Mills system differ substantially from those employed in the study of gravity coupled with massless matter fields [10, 20, 2].
@@
In this paper, we initiate the study of weak cosmic censorship conjecture for the Einstein Yang–Mills system in spherical symmetry. We will first consider the characteristic problem for the spherically symmetric Einstein Yang–Mills (SSEYM) system where the initial data are prescribed on a future light cone emanating from vertex at the center of symmetry. In this setting, we demonstrate the local existence theorem for data with a low-regularity norm ($H^{1}$ norm). We also deduce an extension principle for developments of $H^{1}$ data. Following this extension theorem, the Cauchy stability is justified for the Einstein Yang–Mills system under spherically symmetric perturbations, which are initially small with respect to the $H^{1}$ norms.
As will be discussed later, the method used for the Einstein massless scalar field fails in the Einstein Yang–Mills scenario and hence an extension principle within the BV topology is not expected. The new $L^{2}$-based method we propose in this paper has a closer nature to the vacuum setting, and may be relevant for any future progress in vacuum. Indeed, our method is intimately tied to the Hölder topology that we are currently working in progress.
@@ 1.1. Statement of the main results
We formulate the SSEYM system with the gauge group $SU(2)$ in Section 2.1, where the spherically symmetric purely magnetic ansatz for the Yang–Mills potential (2.2) is assumed. The characteristic problem of the SSEYM system involves prescribing a $C^{0}$ function $\partial_{r}w|_{C_{u_{0}}}$ on an initial outgoing cone $C_{u_{0}}$ that emanates from vertex at $\Gamma$. Here $w$ is related to the Yang–Mills potential (2.2). Throughout this paper, the $C^{0}$ Yang–Mills solution always refers to $\partial_{r}w\in C^{0}$.
@@
This suggests that instead of specifying the initial datum $w|_{C_{u_{0}}}$, one may equivalently prescribe $\partial_{r}w|_{C_{u_{0}}}$.
@@
$\|\partial_{r}w\|_{H^{1}}:=\int_{C_{u}}\frac{1}{r}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}+(r\partial_{r}(\partial_{r}w))^{2}\right)\mathrm{d}r.$ (1.4)
@@
$\int_{0}^{r}\frac{1}{r}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r.$ (1.5)
@@
$\int_{0}^{r_{1}}\frac{1}{r}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r=\int_{-\infty}^{\ln r_{1}}\sum_{0\leq k\leq 2}\left(\partial_{s}^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}s.$ (1.6)
The norm (1.6) is the standard $H^{2}$ norm for $\frac{\hat{w}}{r}$. We abuse the notation a little bit and call (1.6) the $H^{2}$ norm of $\frac{\hat{w}}{r}$, or rather $H^{1}$ norm of $\partial_{r}w$.
@@
As a further remark, the standard energy for the Yang–Mills field does not align with the definition of energy given in (1.4). In fact, the conventional first-order energy for the Yang-Mills field takes the form
$\int_{C_{u}}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}\right)\mathrm{d}r,$
which lacks the weight factor $\frac{1}{r}$. The motivation for (1.4) stems from the right hand side of (2.14a): after integrating (2.14a) in $\underline{u}$, we aim to bound $\frac{m}{r}$ via the $H^{1}$ (in fact, $H^{0}$) norm of $\partial_{r}w$.
We now state the $H^{1}$-extension theorem. A complete and detailed formulation is provided in Theorem 3.1.
@@
Suppose we have initial data $\partial_{r}w|_{C_{u_{0}}}$ with bounded $H^{1}$ norm, and let $w$ be an $H^{1}$ solution (that is, $\partial_{r}w\in H^{1}$) to the SSEYM system on the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$. There is a constant $\varepsilon>0$ such that if
@@
the solution extends as an $H^{1}$ solution on the domain $\mathcal{D}(u_{0};\,\underline{u}_{1})$ for some $\underline{u}_{1}>\underline{u}_{\ast}$.
Following the extension theorem, we state a Cauchy stability result. A precise version is presented in Theorem 3.7.
@@
Suppose the data $\partial_{r}w|_{C_{u_{0}}}$ admit a bounded $H^{1}$ norm that is small enough. Then, for any $\underline{u}_{0}\in(u_{0},+\infty)$ and $u_{1}\in(u_{0},\,\underline{u}_{0}]$, there exists a unique and global solution in $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{0})$ whose $H^{1}$ norm remains small.
@@
In the subsection followed, we will sketch the main idea for the proofs. We introduce the $L^{2}$-based method, tailored to the Yang–Mills field–or more precisely, to wave type equations with a positive, singular potential. This approach can also be adapted to both massless and massive scalar fields, ensuring that all the preceding statements (Theorems 1.41.5) remain valid for the spherically symmetric Einstein (massive or massless) scalar field system.
Since our proof is independ of $C^{1}$ bounds, but only depends on $H^{1}$ estimates, we effectively establish an extension principle in a weaker topology–the $H^{1}$ topology, compared to the $C^{1}$ framework used by Christodoulou [10]. Consequently, our results also require less regularity along the axis.
To facilitate the extension theorems, we establish a local existence theory for the characteristic problem of the SSEYM system in Section 4. Additionally, for completeness, we present extension principle away from the axis in Appendix A. Thus, it confirms that the “first singularity” for the SSEYM system must orginate from the axis.
@@ 1.2. Comments on the proof
The purely magnetic Yang–Mills field obeys a wave-type equation with a singular potential proportional to $\frac{1}{r}$, which distinguishes it fundamentally from a standard massless scalar field. To highlight the difference, it is instructive to compare their formulations in Minkowski spacetime. The massless scalar field equation in Minkowski takes the form
@@
while the Yang–Mills equation is given by
@@
In terms of the variable $\hat{w}$ defined in (1.3), the Yang–Mills equation can be reformulated as
@@
This exhibits a nonlinear wave structure with an additional linear, singular term
@@
absent in the scalar wave case. As a result, the methods required to establish extension principles in the two settings necessarily differ.
@@ 1.2.1. The $H^{1}$ case
Remarkably, the right-hand side of this equation contains no linear terms in $\partial_{r}\alpha_{s}$, $\alpha_{s}$, $r\partial_{r}\phi$ and $\frac{2m}{r}$. It is exactly this structure that enables a usage of an $L^{\infty}-L^{\infty}$ estimate in the proof of the extension theorem [10, Theorem 5.1].
To illustrate the key idea of this method, consider, without loss of generality, the term
$2\frac{\partial_{u}r}{r}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}\partial_{r}\alpha_{s}.$
Its treatment exemplifies how nonlinear contributions are controlled within the $L^{\infty}$ framework. In this quadratic term, the $\frac{2m}{r}$ offers a smallness factor $\varepsilon$, rendering the entire term negligible. This allows an inequality of the form
@@
to yield a bound for $\partial_{r}\alpha_{s}$. In principle, such $L^{\infty}-L^{\infty}$ estimates are effective for “generic” massless matter fields that share the characteristic absence of linear terms. Besides the massless scalar field, this $L^{\infty}-L^{\infty}$ approach has also been shown to apply to the charged massless scalar field [2] and the massless Vlasov system [20], to the best of our knowledge.
In contrast, for the Yang–Mills case, the presence of an additional potential leads to linear terms (linear in $\frac{\hat{w}}{r}$ and $\alpha$) in the main equation for $\partial_{r}\alpha$ (2.3), $\alpha:=\partial_{r}w$. Consequently, the $L^{\infty}-L^{\infty}$ estimate strategy fails here, as these linear terms do not provide the required smallness.
Nevertheless, the positive sign of the potential in the Yang-Mills equation motivates the use of an $L^{2}$-based method. This approach allows us to handle quadratic terms arising from the linear contributions via integration by parts. After this manipulation, terms lacking inherent smallness give rise to quadratic expressions with definite signs. However, since our energy is weighted by $r^{-1}$ (and thus non-standard), these quadratic terms do not automatically exhibit favorable signs. In particular, the first-order energy identities (3.5) and (3.5) do not directly yield an energy bound.
For this issue, we proceed to the second-order energy estimates and integrating by parts on the linear-derived terms. The second-order energy identity then produces quadratic terms with definite signs, which can potentially compensate for corresponding terms of opposite sign in the first-order identities (see (3.5)). By strategically weighting these two energy identities, we ensure that all quadratic spacetime integrals become positive. This ultimately yields $H^{1}$ energy bounds (as detailed in Section 3.5), and pointwise estimates follow via a Sobolev-type inequality in the $H^{1}$ space.
The method developed for the Yang-Mills case can also be applied to both massless and massive scalar fields, enabling the establishment of an extension principle in a low-regularity ($H^{1}$) setting. Furthermore, our approach allows less regularity near the axis: we only require $r\partial_{r}\phi\sim 1$, while in Christodoulou [10], the setting $r\partial_{r}\phi\sim r$ near the axis is stronger. Additionally, the $H^{1}$ energy estimates play a crucial role in establishing the local existence result given in Theorem 4.1.